OpenStudy (thecatman):

how do you simplify 4x+37>=-15?

OpenStudy (texaschic101):

4x + 37 >= -15 --- subtract 37 from both sides 4x >= -15 - 37 4x >= - 52 -- divide both sides by 4 x >= -52/4 x >= - 13

OpenStudy (texaschic101):

you basically solve it just like equations....but with a few rules.....with inequalities, when dividing/multiplying by a negative, the sign changes.

OpenStudy (thecatman):

oops i meant 4x=3y>=-15 sorry

OpenStudy (texaschic101):

I am not sure....sorry

OpenStudy (thecatman):

its ok my teacher can explain it tomorrow

OpenStudy (texaschic101):

4x = 3y >= -15 ---- possibly divide by 3 4/3x = y > = -5 not real sure....

OpenStudy (thecatman):

can it be graphed instead

OpenStudy (texaschic101):

not sure...you might need to check with your teacher...or get somebody else on here that can possibly help

OpenStudy (texaschic101):

sorry :(

OpenStudy (thecatman):

i got this

OpenStudy (thecatman):

as my answer

OpenStudy (thecatman):

is it correct or was i off by alot

OpenStudy (anonymous):

\[4x \ge-15,x \ge \frac{ -15 }{ 4 }or ~x \ge-3.75\] \[3y \ge-15,y \ge -5\]

OpenStudy (anonymous):



OpenStudy (thecatman):

thank you for the help. i owe you one

OpenStudy (anonymous):



OpenStudy (anonymous):

y=4/3 x it is a line passing through origin when x=3,y=4 draw aline through (0,0) and (3,4)